OpenStudy (anonymous):

x2 + y2 + 14x + 18y = -114

OpenStudy (anonymous):

are there two equations or just the one?

OpenStudy (anonymous):

i think its just the one..

OpenStudy (anonymous):

are you perhaps supposed to find the center and radius of this circle?

OpenStudy (anonymous):

\[x^2+14x+y^2+18y=-114\] \[(x+7)^2+(y+9)^2=-114+7^2+9^2=-114+49+81=16\] \[(x+7)^2+(y+9)^2=16\] is the equation for a circle with center at (-7,-9) and radius 4